OpenStudy (anonymous):

Shanti has just joined a DVD rental club. She pays a monthly membership fee of $4.95, and each DVD rental is $1.95. If Shanti's budget for DVD rentals in a month is $42.00, how many DVDs can Shanti rent in her first month if she doesn't want to go over her budget? ________ DVDs

OpenStudy (anonymous):

4.95 + 1.95x = 42 1.95x = 42 - 4.95 1.95x = 37.05 x = 37.05/1.95 x = 19 check... 4.95 + 1.95(19) = 42 4.95 + 37.05 = 42 42 = 42 (correct) She can rent 19 DVD's